OpenStudy (anonymous):

Graph the parametric curve including an arrow to show the direction of the curve: x(t) = t^5 y(t) = -t^2

OpenStudy (anonymous):

@jim_thompson5910 Do you think you could help me?

jimthompson5910 (jim_thompson5910):

First isolate t in x(t) = t^5

jimthompson5910 (jim_thompson5910):

\[\Large x = t^5\] \[\Large \sqrt[5]{x} = t\] \[\Large t=\sqrt[5]{x}\]

jimthompson5910 (jim_thompson5910):

Then plug this into the second equation \[\Large y = -t^2\] \[\Large y = -(\sqrt[5]{x})^2\] \[\Large y = -(x^{1/5})^2\] \[\Large y = -x^{2/5}\]